OpenStudy (kainui):

I'm looking at Grandi's series and trying to figure out what you can't get it to converge to. Grandi's series is: 1-1+1-1+1-... to infinity

OpenStudy (kainui):

So for example you can look at it as: S=(1-1)+(1-1)+...=0 S=1+(-1+1)+(-1+1)+...=1 or something like: S=1-1+1-1+... 1-S=1-1+1-1+... 1-S=S 1=2S S=1/2 But you can do more like: S=(1+1)-(1+1)+(1+1)-...=2 or S=S^2=S^3=S^4... I'm trying to see if I can't make it equal pi, e, or i. Any thoughts or ideas or am I alone on this one haha.

OpenStudy (anonymous):

you could make it equal to \(\pi\) or \(e\) but not \(i\) (I think...); playing with divergent real series (since they have no actual sum) can yield any real number, just like how in logic you can assume inconsistent axioms and derive anything: http://en.wikipedia.org/wiki/Principle_of_explosion